OpenStudy (anonymous):

Use b^2-4ac find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (saifoo.khan):

a = 2 b = 7 c = 3 Now insert the values in: b^2-4ac

OpenStudy (anonymous):

i did but i got 24= 0?

OpenStudy (anonymous):

actually i got 25= 0

OpenStudy (anonymous):

I keep getting diffrent numbers for my answer like73 and 13?

OpenStudy (saifoo.khan):

49 - 4(2)(3) 49 - 24 25

OpenStudy (saifoo.khan):

lol, 25 is the right answer.

OpenStudy (anonymous):

b^2-4ac a=2 b=7 c=3 7^2-4(2)(3)= 49-24= 25

OpenStudy (anonymous):

lol cool so i got it right the first time does it matter that 25=0?

OpenStudy (saifoo.khan):

Lol, congrats.. yes it does.

OpenStudy (anonymous):

wait what do u mean 25=0?

OpenStudy (saifoo.khan):

u dont have to write the answer that way..

OpenStudy (saifoo.khan):

like 25 = 0

OpenStudy (anonymous):

With the equation , but the problem is still correct?

OpenStudy (anonymous):

no 25 does not equal zero. the formula is \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (saifoo.khan):

the answer will be just 25.

OpenStudy (anonymous):

no the answer will not be 25

OpenStudy (saifoo.khan):

o.O

OpenStudy (anonymous):

the answer will be \[\frac{-7\pm\sqrt{25}}{4}\]

OpenStudy (saifoo.khan):

she wants the discriminent.

OpenStudy (anonymous):

he's right the answer isnt 25 cuz u ddnt use the whole quadratic equation

OpenStudy (anonymous):

no it says solve using discriminant

OpenStudy (anonymous):

no she wants the answer which is y its set equal to 0 but she ddnt put the whole formula down

OpenStudy (saifoo.khan):

-_- Confused.

OpenStudy (anonymous):

u have to use the discriminant to find the x values that solve the equation like satellite said

OpenStudy (anonymous):

\[b^2-4ac=25\] and the solutions to \[2x^2 + 7x + 3 = 0\]using that in the formula are \[\{-\frac{1}{2},-3\}\]

OpenStudy (saifoo.khan):

in other words we have to use quadratic formula!

OpenStudy (anonymous):

yup but written in a confusing way

OpenStudy (saifoo.khan):

Leave it..

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

....okay all of that was a cunfusing way of telling me that 25 is correct????

OpenStudy (saifoo.khan):

i will go with 25, rest go with satellite's answer.

OpenStudy (anonymous):

the answer the way u aske it is 25 because u asked for the discriminant. but what exactly did u mean 25=0?

OpenStudy (anonymous):

just i didnt know if you used the equation at the top or not